# Efficient multi-prompt evaluation of LLMs

### v2

@@ Abstract
Most popular benchmarks for comparing LLMs rely on a limited set of prompt templates, which may not fully capture the LLMs’ abilities and can affect the reproducibility of results on leaderboards. Many recent works empirically verify prompt sensitivity and advocate for changes in LLM evaluation. In this paper, we consider the problem of estimating the performance distribution across many prompt variants instead of finding a single prompt to evaluate with. We introduce PromptEval, a method for estimating performance across a large set of prompts borrowing strength across prompts and examples to produce accurate estimates under practical evaluation budgets. The resulting distribution can be used to obtain performance quantiles to construct various robust performance metrics (e.g., top 95% quantile or median). We prove that PromptEval consistently estimates the performance distribution and demonstrate its efficacy empirically on three prominent LLM benchmarks: MMLU, BIG-bench Hard, and LMentry. For example, PromptEval can accurately estimate performance quantiles across 100 prompt templates on MMLU with a budget equivalent to two single-prompt evaluations [^1] Our code can be found in https://github.com/felipemaiapolo/prompteval. The MMLU data can be found in https://huggingface.co/PromptEval..
@@ 1 Introduction
We introduce PromptEval, a method for efficient multi-prompt evaluation of LLMs. With a small number of evaluations, PromptEval estimates performance across different prompt templates. Our approach is grounded in robust theoretical foundations and utilizes well-established models from the fields of educational assessment and psychometrics, such as Item Response Theory (IRT) (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968). Our method is based on a parametric IRT model that allows borrowing strength across examples and prompt templates to produce accurate estimates of all considered prompts with an evaluation budget comparable to evaluating a single prompt. In Figure 1, we demonstrate the ability of our method to jointly estimate various performance quantiles across 100 prompt templates with evaluation budget ranging from one to four times of a conventional single-prompt evaluation on MMLU (Hendrycks et al., 2020).
Performance distribution across prompts can be used to accommodate various contexts when comparing LLMs (Choshen et al., 2024). For example, it can be used to compute mean as suggested by Mizrahi et al. (2023). One can also use performance distributions directly to compare LLMs via various notions of stochastic dominance for risk-sensitive scenarios (Nitsure et al., 2023). Here we primarily focus on the full distribution and its quantiles as they provide a flexible statistic that can inform decisions in varying contexts. For instance, a typical model performance corresponds to a median (50% quantile), 95% quantile can be interpreted as performance achievable by an expert prompt engineer, while 5% quantile is of interest in consumer-facing applications to quantify low-end performance for a user not familiar with prompt engineering. We also demonstrate (§5) how our method can be used to improve best prompt identification (Shi et al., 2024).
@@
  • We propose (§3) a novel method called PromptEval which permits efficient multi-prompt evaluation of LLMs across different prompt templates with a limited number of evaluations.
  • We theoretically show (§4) that PromptEval has desirable statistical properties such as consistency in estimating performance distribution and its quantiles.
@@
  • We conduct the first large-scale study of prompt sensitivity of 15 popular open-source LLMs on MMLU. We present our findings based on evaluating 100 prompt templates in Section 6 and will release the evaluation data.
@@ 1.1 Related work
Efficient evaluation of LLMs The escalating size of models and datasets has led to increased evaluation costs. To streamline evaluations, Ye et al. (2023b) considered minimizing the number of tasks within Big-bench (Srivastava et al., 2022). Additionally, Perlitz et al. (2023) observed that evaluations on HELM (Liang et al., 2022) rely on diversity across datasets, though the quantity of examples currently utilized is unnecessarily large. Perlitz et al. (2023) also highlighted the problems in evaluating with insufficient prompts and called to evaluate on more, suggesting evaluating the typical behavior by sampling prompts and examples together by employing stratified sampling, where subscenarios give the strata; in our work, we also apply stratification but consider prompt templates and examples to give the strata. To accelerate evaluations for classification tasks, Vivek et al. (2023) suggested clustering evaluation examples based on model confidence in the correct class. More recently, Polo et al. (2024) empirically showed that it is possible to shrink the size of modern LLM benchmarks and still retain good estimates for LLMs’ performances. Similarly (and in parallel to this work) Ashury-Tahan et al. (2024) recognized unlabeled examples that better distinguish between models or prompts, by analyzing model outputs on them, hence saving costly annotation for them. Despite these advancements in streamlining LLM evaluations, there are no other works that propose a general and efficient method to estimate the benchmark performance of LLMs across prompt templates to the best of our knowledge.
Item response theory (IRT) IRT (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968) is a collection of statistical models initially developed in psychometrics to assess individuals’ latent abilities through standardized tests but with increasing importance in the fields of artificial intelligence and natural language processing (NLP). For example, Lalor et al. (2016) used IRT’s latent variables to measure language model abilities, Vania et al. (2021) applied IRT to benchmark language models and examine the saturation of benchmarks, and Rodriguez et al. (2021) explored various uses of IRT with language models, including predicting responses to unseen items, categorizing items by difficulty, and ranking models. Recently, Polo et al. (2024) suggested using IRT for efficient LLM performance evaluation, introducing the Performance-IRT (pIRT) estimator to evaluate LLMs. Our quantile estimation methodology is built upon pIRT.
@@ 3.1 The correctness model
The functions $f_{\psi}$ and $g_{\gamma}$ can be represented with neural networks. On the simpler side, one could just assume $f_{\psi}$ and $g_{\gamma}$ are linear, that is, $\theta_{i}=\psi^{\top}x_{i}$ or $\beta_{j}=\gamma^{\top}z_{j}$. We consider that, in some cases, a constant can be embedded in $x_{i}$ in order to include an intercept in the model. When $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros but with 1’s on the entries $i$ and $j$, the model in (3.2) reverts to a popular IRT model known as the Rasch model (Georg, 1960; Chen et al., 2023), which is widely used in fields such as recommendation systems (Starke et al., 2017) and educational testing (Clements et al., 2008). One major limitation of the basic Rasch model is that the number of parameters is large, compromising the quality of the estimates for $\psi$ and $\gamma$ when either the number of prompt formats $I$ or the number of examples $J$ is large and $|\mathcal{E}|$ is small, i.e., only a few evaluations are carried out. This degradation in the quality of the estimates can directly affect the quality of the performance distribution estimates. Finally, we fit the parameters $\psi$ and $\gamma$, obtaining the estimates $\hat{\psi}$ and $\hat{\gamma}$, by maximizing the log-likelihood of the observed data (negative cross-entropy loss), i.e.,
@@
Realize that fitting the model with linear/affine $f_{\psi}$ and $g_{\gamma}$, including the Rasch model case [^3] For a detailed fitting procedure in the Rasch model case, please check Chen et al. (2023)., reduces to fitting a logistic regression model with $x_{i}$ and $z_{j}$ as the covariates. In the experiments Section 5, we explore some different options of covariates for both templates and examples.
@@ 3.2 Performance distribution and quantiles estimation using the correctness model
The basic version of this estimator, when no elaborate covariates (e.g., embeddings) are included, is known as the Performance-IRT (pIRT) estimator (Polo et al., 2024). We can apply our extended version of pIRT, which we call X-pIRT, to estimate the performance distribution across prompt templates. After observing $Y_{\mathcal{E}}$ and fitting $(\hat{\psi},\hat{\gamma})$, we can compute $\hat{S}_{i}\triangleq\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ for all $i\in\mathcal{I}$. Then, we define our estimators for the distribution of performances and its corresponding quantile function 2.1 as
@@ 4 Theoretical guarantees
In this section, we claim the consistency of the distribution and quantile estimators detailed in Algorithm 1 as $I,J\to\infty$. We prove a result for the case in which $f_{\psi}$ and $g_{\gamma}$ are linear/affine functions. Before we introduce our results we need to introduce some basic conditions. As an extra result, in Appendix F.1 we also show that our extended version of pIRT (3.4), X-pIRT, is uniformly consistent over all $i\in\mathcal{I}$, which can be useful beyond this work.
We start by assuming that the covariates are uniformly bounded.
@@
We now introduce the main result in Theorem 4.4, which shows the consistency of the distribution and quantile functions estimators introduced in Algorithm 1. See Appendix F for the proof.
@@ 5 Assessing multi-prompt evaluation strategies
  • MMLU is a multiple choice QA benchmark consisting of 57 subjects (tasks) comprising approximately 14k examples. We ran 15 different open-source LLMs (including different versions of Llama-3 (Meta, 2024), Mistral (Jiang et al., 2023), and Gemma (Gemma et al., 2024)) combined with 100 different prompt variations for each one of the MMLU tasks. We found that, within each one of the MMLU tasks, prompt templates can have great variability in their performances, making within-task analysis most suitable for assessing our method. More details and analysis of the collected data can be found in §6 and Appendix G.
@@
We consider different variations of the model presented in (3.2) coupled with Algorithm 1; for all variations, we use linear $f_{\psi}$ and $g_{\gamma}$. The most basic version of the model in (3.2) assumes $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros with 1’s on the entries $i$ and $j$, reverting the model to a Rasch model (Georg, 1960; Chen et al., 2023). Despite its simplicity, we show that it can perform well in some cases. A more advanced instance of (3.2) assumes $x_{i}$ are either obtained using a sentence transformer (Reimers and Gurevych, 2019) to embed prompt templates or by extracting discrete covariates from the text, e.g., as the presence of line breaks, colons etc.(see Appendix Table 2). An example of a prompt template for LMentry used by Mizrahi et al. (2023) is “Can {category} be used to classify all the {words} provided? Respond with either "yes" or "no".” Our method also allows using example covariates $z_{j}$, however, upon preliminary tests with sentence transformer we didn’t observe improvements and chose to use one-hot-encoded vectors as in the basic Rasch model to represent examples. Next, we detail the methods for obtaining the prompt covariates:
@@
  • Fine-tuned prompt embeddings. Sentence transformers in general might not be most suitable for embedding prompt templates, thus we also consider fine-tuning BERT (Devlin et al., 2019) as an embedder. To do so, we use evaluation data for all examples and prompt formats from a subset of LLMs (these LLMs are excluded when assessing the quality of our estimators) and fine-tune bert-base-uncased to predict $Y_{ij}$ as in (3.3). We call this variation EmbFT and provide additional details in Appendix I. We acknowledge that obtaining such evaluation data for fine-tuning might be expensive, however, it might be justified in some applications if these embeddings provide sufficient savings for future LLM evaluations.
  • Discrete prompt covariates. For BBH and LMentry, we coded a heuristic function that captures frequently occurring differences in common prompting templates. Examples of such covariates are the number of line breaks or the count of certain special characters (e.g., dashes or colons). Each one of these covariates is encoded in $x_{i}$ for each one of the prompt templates $i\in\mathcal{I}$. A full list of the used heuristics is detailed in Appendix J. For MMLU, we adopted approach of (Sclar et al., 2023) to generate prompt variations via templates (see Algorithm 3), which also provides a natural way to construct the covariates, e.g., the presence of dashes or colons.
@@
Best-prompt identification

### v3

@@ Abstract
Most popular benchmarks for comparing LLMs rely on a limited set of prompt templates, which may not fully capture the LLMs’ abilities and can affect the reproducibility of results on leaderboards. Many recent works empirically verify prompt sensitivity and advocate for changes in LLM evaluation. In this paper, we consider the problem of estimating the performance distribution across many prompt variants instead of finding a single prompt to evaluate with. We introduce PromptEval, a method for estimating performance across a large set of prompts borrowing strength across prompts and examples to produce accurate estimates under practical evaluation budgets. The resulting distribution can be used to obtain performance quantiles to construct various robust performance metrics (e.g., top 95% quantile or median). We prove that PromptEval consistently estimates the performance distribution and demonstrate its efficacy empirically on three prominent LLM benchmarks: MMLU, BIG-bench Hard, and LMentry; for example, PromptEval can accurately estimate performance quantiles across 100 prompt templates on MMLU with a budget equivalent to two single-prompt evaluations. Moreover, we show how PromptEval can be useful in LLM-as-a-judge and best prompt identification applications. [^1] Our code can be found in https://github.com/felipemaiapolo/prompteval and the MMLU data can be found in https://huggingface.co/PromptEval. PromptEval is also integrated into PromptBench (Zhu et al., 2024); please check https://github.com/microsoft/promptbench.
@@ 1 Introduction
We introduce PromptEval, a method for efficient multi-prompt evaluation of LLMs. With a small number of evaluations, PromptEval estimates performance across a large and given pool of different prompt templates. Our approach is grounded in robust theoretical foundations and utilizes well-established models from the fields of educational assessment and psychometrics, such as Item Response Theory (IRT) (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968). Our method is based on an IRT model that allows borrowing strength across examples and prompt templates to produce accurate estimates of all considered prompts with an evaluation budget comparable to evaluating a single prompt. In Figure 1, we demonstrate the ability of our method to jointly estimate various performance quantiles across 100 prompt templates with an evaluation budget ranging from one to four times of a conventional single-prompt evaluation on MMLU (Hendrycks et al., 2020).
Performance distribution across prompts can be used to accommodate various contexts when comparing LLMs (Choshen et al., 2024). For example, it can be used to compute the mean performance as suggested by Mizrahi et al. (2023). One can also use performance distributions directly to compare LLMs via various notions of stochastic dominance for risk-sensitive scenarios (Nitsure et al., 2023). Here we primarily focus on the full distribution and its quantiles as they provide a flexible statistic that can inform decisions in varying contexts. For instance, a typical model performance corresponds to a median (50% quantile), 95% quantile can be interpreted as performance achievable by an expert prompt engineer, while 5% quantile is of interest in consumer-facing applications to quantify low-end performance for a user not familiar with prompt engineering. We also demonstrate (§6) how our method can be used to account for prompt sensitivity in the LLM-as-a-judge framework (Li et al., 2023) and to do best prompt identification (Shi et al., 2024).
@@
  • We propose (§3) a novel method called PromptEval which permits efficient multi-prompt evaluation of LLMs for a given pool of prompt templates with a limited number of evaluations. Moreover, we theoretically show (§4) that PromptEval has desirable statistical properties such as consistency in estimating performance distribution and its quantiles.
@@
  • We show (§6) how PromptEval can be applied to account for prompt sensitivity in the LLM-as-a-judge framework and to identify the best prompt in a large pool of options.
  • We conduct the first large-scale study of prompt sensitivity of 15 popular open-source LLMs on MMLU. We present our findings based on evaluating 100 prompt templates in Section 7 and release the evaluation data.
@@ 1.1 Related work
Efficient evaluation of LLMs The escalating size of models and datasets has led to increased evaluation costs. To streamline evaluations, Ye et al. (2023b) considered minimizing the number of tasks within Big-bench (Srivastava et al., 2022). Additionally, Perlitz et al. (2023) observed that evaluations on HELM (Liang et al., 2022) rely on diversity across datasets, though the quantity of examples currently utilized is unnecessarily large. Perlitz et al. (2023) also highlighted the problems in evaluating with insufficient prompts and called to evaluate on more, suggesting evaluating the typical behavior by sampling prompts and examples together by employing stratified sampling, where subscenarios give the strata; in our work, we also apply stratification but consider prompt templates and examples to give the strata. To accelerate evaluations for classification tasks, Vivek et al. (2023) suggested clustering evaluation examples based on model confidence in the correct class. More recently, Maia Polo et al. (2024) empirically showed that it is possible to shrink the size of modern LLM benchmarks and still retain good estimates for LLMs’ performances. Similarly (and in parallel to this work) Ashury-Tahan et al. (2024) recognized unlabeled examples that better distinguish between models or prompts, by analyzing model outputs on them, hence saving costly annotation for them. Despite these advancements in streamlining LLM evaluations, there are no other works that propose a general and efficient method to estimate the benchmark performance of LLMs across prompt templates to the best of our knowledge.
Item response theory (IRT) IRT (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968) is a collection of statistical models initially developed in psychometrics to assess individuals’ latent abilities through standardized tests but with increasing importance in the fields of artificial intelligence and natural language processing (NLP). For example, Lalor et al. (2016) used IRT’s latent variables to measure language model abilities, Vania et al. (2021) applied IRT to benchmark language models and examine the saturation of benchmarks, and Rodriguez et al. (2021) explored various uses of IRT with language models, including predicting responses to unseen items, categorizing items by difficulty, and ranking models. Recently, Maia Polo et al. (2024); Shabtay et al. (2024) suggested using IRT for efficient LLM performance evaluation; both works used the Performance-IRT (pIRT) estimator to evaluate LLMs. PromptEval is built upon pIRT.
@@ 3.1 The correctness model
The functions $f_{\psi}$ and $g_{\gamma}$ can be represented with neural networks. On the simpler side, one could just assume $f_{\psi}$ and $g_{\gamma}$ are linear, that is, $\theta_{i}=\psi^{\top}x_{i}$ or $\beta_{j}=\gamma^{\top}z_{j}$; this formulation is known as the linear logistic test model in psychometrics (Fischer, 1973; De Boeck, 2004). We consider that, in some cases, a constant can be embedded in $x_{i}$ in order to include an intercept in the model. When $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros but with 1’s on the entries $i$ and $j$, the model in (3.2) reverts to a popular IRT model known as the Rasch model (Georg, 1960; Chen et al., 2023), which is widely used in fields such as recommendation systems (Starke et al., 2017) and educational testing (Clements et al., 2008). One major limitation of the basic Rasch model is that the number of parameters is large, compromising the quality of the estimates for $\psi$ and $\gamma$ when either the number of prompt formats $I$ or the number of examples $J$ is large and $|\mathcal{E}|$ is small, i.e., only a few evaluations are carried out. This degradation in the quality of the estimates can directly affect the quality of the performance distribution estimates. Finally, we fit the parameters $\psi$ and $\gamma$, obtaining the estimates $\hat{\psi}$ and $\hat{\gamma}$, by maximizing the log-likelihood of the observed data (negative cross-entropy loss), i.e.,
@@
Realize that fitting the model with linear/affine $f_{\psi}$ and $g_{\gamma}$, including the Rasch model case [^3] For a detailed fitting procedure in the Rasch model case, please check Chen et al. (2023)., reduces to fitting a logistic regression model with $x_{i}$ and $z_{j}$ as the covariates. This observation highlights that the fitting process is expected to be very cheap in practice. For example, in our experiments, we fit logistic regression models in datasets with less than 2k samples and a couple of hundred (or a few thousand) columns, which is performed in a few seconds by a modern laptop. We include some more comments on the computational complexity of our method in Appendix C.
@@ 3.2 Performance distribution and quantiles estimation using the correctness model
The basic version of this estimator, when no elaborate covariates (e.g., embeddings) are included, is known as the Performance-IRT (pIRT) estimator (Maia Polo et al., 2024). We can apply our extended version of pIRT, which we call X-pIRT, to estimate the performance distribution across prompt templates. After observing $Y_{\mathcal{E}}$ and fitting $(\hat{\psi},\hat{\gamma})$, we can compute $\hat{S}_{i}\triangleq\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ for all $i\in\mathcal{I}$. Then, we define our estimators for the distribution of performances and its corresponding quantile function 2.1 as
@@ 4 Theoretical guarantees
In this section, we claim the consistency of the distribution and quantile estimators detailed in Algorithm 1 as $I,J\to\infty$. We prove a result for the case in which $f_{\psi}$ and $g_{\gamma}$ are linear/affine functions. Before we introduce our results we need to introduce some basic conditions. As an extra result, in Appendix I.1 we also show that our extended version of pIRT (3.4), X-pIRT, is uniformly consistent over all $i\in\mathcal{I}$, which can be useful beyond this work. We start by assuming that the covariates are uniformly bounded.
@@
We now introduce the main result in Theorem 4.4, which shows the consistency of the distribution and quantile functions estimators introduced in Algorithm 1. See Appendix I for the proof.
@@ 5 Assessing multi-prompt evaluation strategies
  • MMLU is a multiple choice QA benchmark consisting of 57 subjects (tasks) comprising approximately 14k examples. We ran 15 different open-source LLMs (including different versions of Llama-3 (Meta, 2024), Mistral (Jiang et al., 2023), and Gemma (Gemma et al., 2024)) combined with 100 different prompt variations for each one of the MMLU tasks. We found that, within each one of the MMLU tasks, prompt templates can have great variability in their performances, making within-task analysis most suitable for assessing our method. More details and analysis of the collected data can be found in §7 and Appendix J.
@@
We consider different variations of the model presented in (3.2) coupled with Algorithm 1; for all variations, we use linear $f_{\psi}$ and $g_{\gamma}$. The most basic version of the model in (3.2) assumes $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros with 1’s on the entries $i$ and $j$, reverting the model to a Rasch model (Georg, 1960; Chen et al., 2023). Despite its simplicity, we show that it can perform well in some cases. A more advanced instance of (3.2) assumes $x_{i}$ are either obtained using a sentence transformer (Reimers and Gurevych, 2019) to embed prompt templates or by extracting discrete covariates from the text, e.g., as the presence of line breaks, colons etc.(see Appendix Table 2). An example of a prompt template for LMentry used by Mizrahi et al. (2023) is “Can {category} be used to classify all the {words} provided? Respond with either "yes" or "no".” Our method also allows using example covariates $z_{j}$, however, upon preliminary tests with sentence transformer we didn’t observe improvements and chose to use one-hot-encoded vectors as in the basic Rasch model to represent examples. Next we detail the methods for obtaining the prompt covariates:
@@
  • Fine-tuned prompt embeddings. Sentence transformers in general might not be most suitable for embedding prompt templates, thus we also consider fine-tuning BERT (Devlin et al., 2019) as an embedder. To do so, we use evaluation data for all examples and prompt formats from a subset of LLMs (these LLMs are excluded when assessing the quality of our estimators) and fine-tune bert-base-uncased to predict $Y_{ij}$ as in (3.3). We call this variation EmbFT and provide additional details in Appendix L. We acknowledge that obtaining such evaluation data for fine-tuning might be expensive, however, it might be justified in some applications if these embeddings provide sufficient savings for future LLM evaluations.
  • Discrete prompt covariates. For BBH and LMentry, we coded a heuristic function that captures frequently occurring differences in common prompting templates. Examples of such covariates are the number of line breaks or the count of certain special characters (e.g., dashes or colons). Each one of these covariates is encoded in $x_{i}$ for each one of the prompt templates $i\in\mathcal{I}$. A full list of the used heuristics is detailed in Appendix M. For MMLU, we adopted approach of (Sclar et al., 2023) to generate prompt variations via templates (see Algorithm 3), which also provides a natural way to construct the covariates, e.g., the presence of dashes or colons.
@@
6 Further applications of PromptEval
6.1 Estimating the distribution of scores for the LLM-as-a-judge framework
In this subsection, we explore the concept of LLM-as-a-judge using the AlpacaEval 2.0 (Li et al., 2023) benchmark. Specifically, we generate 100 prompt templates [^4] We generate 10k variations using ChatGPT and undersample to 100 deleting prompts that are too similar to each other. You can see our code here. to present to the judge, GPT-4o-mini (OpenAI, 2024), allowing us to assess how sensitive model evaluation is to different evaluation prompts. We evaluated the performance of four LLMs with similar capabilities (Cohere Command [^5] https://docs.cohere.com/v2/docs/command-beta, Qwen1.5-7B-Chat (Team, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), LLaMa-2-70B-Chat (Touvron et al., 2023)) using only $\approx 2\%$ of the total evaluations (1.6k/80.5k). In contrast to the previous experiments, we do not make changes in the prompt templates given to the evaluated LLMs when giving an instruction. To fit PromptEval, we binarize AlpacaEval 2.0 instance scores imposing a threshold at $1/2$, but we do not binarize the responses at test time. Figure 4 shows that the different variations of PromptEval can obtain a much lower Wasserstein loss ($W_{1}$) when compared with the baseline “avg”. In Appendix G, we provide additional plots for this experiment. Specifically, Figure 11 presents the score distribution histograms for the four models under consideration, while Figure 12 illustrates how certain prompt templates consistently lead the judge to assign higher (or lower) scores across models. Despite this pattern, we observe that the ranking of the four LLMs changes in 36% of the prompt templates.
Figure 4: Estimating LLM-as-a-judge distribution of scores for 100 prompt variations given to the judge.
6.2 Best-prompt identification